OpenStudy (anonymous):

Find the area of a regular hexagon with an apothem of 9 in. and a side length of 12 in. A. 432 sq. in. B. 324 sq. in. C. 243 sq. in. D. 54 sq. in.

OpenStudy (anonymous):

approach the problem by splitting it into triangles. use the apothem as the height.

OpenStudy (anonymous):

Area = 1/2 * (perimeter) * (apothem) Area = 1/2 * (6 * sides) * (apothem)

OpenStudy (anonymous):

well if you have the formulas then just plug and chug lol... 1/2*(12*6)*12

OpenStudy (anonymous):

1/2*(12*6)*9 = 324? which is B ! thanks @sylbot

OpenStudy (anonymous):

oops I said 12, I meant 9, youre right,